OpenStudy (anonymous):

Anyone know how to solve this? (2y)-3= ????

OpenStudy (anonymous):

Incomplete info, re-read question

OpenStudy (anonymous):

\[\implies 2y = ???? + 3\]\[\implies y = {???? + 3 \over 2}\]

OpenStudy (anonymous):

Sorry! x=6, y=5, z=1/2 this is evaluating expressions.

OpenStudy (anonymous):

For 'evaluating' you just plug in the value you are given for the variable.

OpenStudy (anonymous):

i guess ur ???? will in most general sense be 0 so (2y)-3=0 will give y=3/2 or as polpak said y= (?????+3)/2

OpenStudy (anonymous):

So \[2y - 3\]\[=2(5)-3\]\[ = 10-3\] ... etc

OpenStudy (anonymous):

Oh! Got it. Thanks!

OpenStudy (anonymous):

But 10 minus 3 is seven. Why would the answer be zero?

OpenStudy (anonymous):

It's not.

OpenStudy (anonymous):

...................

OpenStudy (anonymous):

So where did the zero come from in the beginning?

OpenStudy (anonymous):

What 0?

OpenStudy (anonymous):

The K dude said something about my '?????' being zero

OpenStudy (anonymous):

I have no idea. In this case your ??? was a 7

OpenStudy (anonymous):

i meant most equations take the form of say 2y-3=0 and definitely not 2y-3 =??????

OpenStudy (anonymous):

.........Okay. Thanks

OpenStudy (anonymous):

@satellite73: What?

OpenStudy (anonymous):

y= 3/2